# Bigger, Regularized, Optimistic: scaling for compute and sample-efficient continuous control

### v1

@@ 5 Limitations and Future Work
BRO’s larger model size compared to traditional baselines like SAC or TD3 results in higher memory requirements, potentially posing challenges for real-time inference in high-frequency control tasks. Future research could explore techniques such as quantization or distillation to improve inference speed. While BRO is designed for continuous control problems, its effectiveness in discrete settings remains unexplored.
Further investigation is needed to assess the applicability and performance of BRO’s components in discrete action MDPs. Additionally, our experimentation primarily focuses on continuous control tasks using proprioceptive state representations. Future research could investigate the tradeoff between scaling the critic and the state encoder in image-based continuous control scenarios.
@@ 6 Conclusions
Our study underscores the efficacy of scaling a regularized critic model in conjunction with existing algorithmic enhancements, resulting in sample-efficient methods for continuous-action RL. The proposed BRO algorithm achieves markedly superior performance within 1 million environment steps compared to the state-of-the-art model-based TD-MPC2 and other model-free baselines. Notably, it achieves over 90% success rates in MetaWorld and MyoSuite benchmarks, as well as over 85% of maximal returns in the DeepMind Control Suite, and near-optimal policies in the challenging Dog and Humanoid locomotion tasks.
While some tasks remain unsolved within 1 million environment steps, our findings underscore the need for new standardized benchmarks focusing on sample efficiency to drive consistent progress in the field. The BRO algorithm establishes a new standard for sample efficiency, providing a solid foundation for future research to build upon and develop even more robust RL algorithms.
@@ Appendix A Background
with $a^{\prime}\sim\pi_{\phi}$. In this setup, $Q_{\theta}$ is the critic network, $Q_{\bar{\theta}}$ is the value target network, and $\mathcal{D}$ is the replay buffer (Mnih et al., 2015). $Q_{\theta}$ is trained to approximate the Q-value under the policy from which the bootstrap is sampled (Van Seijen et al., 2009; Sutton & Barto, 2018). The policy parameters $\phi$ are updated to seek locally optimal values approximated by the critic (Ciosek & Whiteson, 2020) according to
@@
$\phi=\arg\max_{\phi}~{}\mathrm{E}_{s\sim\mathcal{D}}~{}\left(Q_{\theta}(s,a)-\alpha\log\pi_{\phi}(a|s)\right)$ (6)
where $a\sim\pi_{\phi}$.
Appendix B Experimental details
B.1 BRO details
B.1.1 Base agent
@@
BRO uses the well-established Soft Actor-Critic (SAC) (Haarnoja et al., 2018) as its base. SAC is a stochastic policy, maximum entropy algorithm (see Eq. 1) that employs online entropy temperature adjustment and an ensemble of two critic networks. SAC models the policy via a Tanh-transformed Gaussian distribution whose parameters are modeled by the actor-network.
B.1.2 Architecture details
@@
In the proposed architecture, we adopt the transformer feedforward blueprint from Vaswani et al. (2017) with a novel layer normalization configuration, as shown in Figure 4. Dropout is omitted. All Dense layers in the BRO have a default width of $512$ units, with a linear layer at both the input and output stages. To increase the model’s depth, we add new residual blocks exclusively. While a similar transformer backbone has been used in previous work (Bjorck et al., 2021), our design choices, detailed in Section C.2, led us to use layer normalization instead of spectral normalization.
B.1.3 Scaling
@@
In Figure 3, we examine the scaling capabilities of SAC and BRO agents using a vanilla dense network (Fujimoto et al., 2018; Haarnoja et al., 2018; Moskovitz et al., 2021), a spectral normalization ResNet (Bjorck et al., 2021; Cetin & Celiktutan, 2023), and a layer normalization ResNet inspired by previous work (Nauman et al., 2024). As shown in Figure 3, increasing the model capacity of a vanilla-dense agent can lead to performance degradation beyond a certain model size. However, for the regularized architectures, we observe behavior similar to the empirically observed scaling properties of supervised models, where increasing model size leads to diminishing returns in performance improvements. Furthermore, we find that the layer normalization ResNet achieves better-scaling properties than the spectral normalization architecture. Interestingly, the BRO agent consistently outperforms the baseline SAC across all architectures and network sizes, suggesting an interaction between parameter scaling and other algorithmic design choices. The highest performing SAC agent achieves around $25\%$ of maximal performance, whereas our proposed BRO agent achieves more than $90\%$. Given that the BRO agent performs similarly at $4.92$ million and $26.31$ million parameters, we use the smaller model to reduce the computational burden.
B.1.4 Scaling replay ratio
@@
B.1.5 Batch Size
Inspired by recent findings that reducing the batch size can result in significant performance gains for discrete action RL (Obando Ceron et al., 2024), we reduce the number of transitions used for gradient calculation from $256$ to $128$. As shown in Figures 8 & 10, this batch size reduction leads to a marginal improvement in aggregate performance and decreased memory requirements of the algorithm. Interestingly, we find that batch size significantly affects performance, with no single value performing best across all tasks.
B.1.6 Risk-Neutral Temporal Difference
@@
Recent studies have shown that techniques like layer normalization or full-parameter resets can be more effective at combating overestimation than pessimistic Q-value approximation (Nauman et al., 2024). Since our critic architecture leverages multiple regularization techniques, we disable CDQ and use the ensemble mean instead of the minimum to calculate the targets for actor and critic updates. In Figures 6 & 8, we compare the performance of the baseline BRO to a BRO agent that uses CDQ. Our findings indicate that using risk-neutral Q-value approximation in the presence of network regularization unlocks significant performance improvements without increasing value overestimation.
B.1.7 Optimistic Exploration
Optimism is an algorithmic design principle that balances exploration and exploitation (Ciosek et al., 2019; Moskovitz et al., 2021). The dual actor setup (Nauman & Cygan, 2023) employs separate policies for exploration and temporal difference updates, with the exploration policy pursuing an optimistic upper-bound Q-value approximation. This approach has been shown to improve the sample efficiency of SAC-based agents (Nauman & Cygan, 2023). We implement the optimistic policy such that the Q-value upper bound is calculated based on the epistemic uncertainty estimated via the quantile critic ensemble (Moskovitz et al., 2021). Figure 6 shows that using a dual policy setup leads to performance improvements. We observe that these results are particularly pronounced in the early training stages and for smaller networks (Figure 8).
@@ B.1.7 Optimistic Exploration
7:     Calculate critic target value without CDQ $Q^{\mu}_{\bar{\theta}}(s^{\prime},a^{\prime})=\frac{1}{2K}\sum_{i=1}^{K}(Q^{k}_{\bar{\theta},1}(s^{\prime},a^{\prime})+Q^{k}_{\bar{\theta},2}(s^{\prime},a^{\prime}))~{}~{}~{}~{}\text{with}~{}~{}~{}~{}a^{\prime}\sim\pi^{p}_{\phi}(s^{\prime})$
8:     Update critic using pessimistic actor $\theta\leftarrow\theta-\nabla_{\theta}\bigl{(}Q_{\theta}(s,a)-(r+\gamma Q^{\mu}_{\bar{\theta}}(s^{\prime},a^{\prime})-\alpha\log\pi^{p}_{\phi}(a^{\prime}|s^{\prime})\bigr{)}^{2}$
9:     Calculate pessimistic actor value without CDQL $Q^{\mu}_{\theta}(s,a)=\frac{1}{2K}\sum_{i=1}^{K}(Q^{k}_{\theta,1}(s,a)+Q^{k}_{\theta,2}(s,a))~{}~{}~{}~{}\text{with}~{}~{}~{}~{}a\sim\pi^{p}_{\phi}(s)$
@@
Algorithm 2 Pseudo-code shows BRO training step, where changes with respect to standard SAC are colored red.
B.2 Tested Environments
We tested BRO on a variety of $40$ tasks from DeepMind Control Suite (Tassa et al., 2018), MyoSuite (Caggiano et al., 2022) and MetaWorld (Yu et al., 2020). Selected tasks cover various challenges, from simple to hard, in locomotion and manipulation. Table 2 presents the environments with specified dimensions of states and actions. BRO is a versatile agent that can successfully perform tasks of different difficulty and various action and state spaces. Our selection of $40$ tasks focuses on the most challenging tasks from the DeepMind Control Suite (DMC), MetaWorld (MW), and MyoSuite (MS) benchmarks, as identified in previous studies (Hansen et al., 2023). We chose these hard tasks because many tasks from these benchmarks can be solved by modern algorithms within 100k environment steps (Hansen et al., 2023; Wang et al., 2024). To address this, previous works have often combined tasks from different benchmarks arbitrarily (Ji et al., 2023; Hansen et al., 2023; Nauman & Cygan, 2023), leading to a lack of standardized benchmarks and slowing progress in the field.
@@ B.2 Tested Environments
Task Observation dimension Action dimension
DeepMind Control
Acrobot-Swingup $6$ $1$
Dog-Trot $223$ $38$
Hopper-Hop $15$ $4$
Humanoid-Run $67$ $24$
Humanoid-Walk $67$ $24$
MetaWorld
Assembly $39$ $4$
Coffee-Push $39$ $4$
Hand-Insert $39$ $4$
Push $39$ $4$
Stick-Pull $39$ $4$
B.3 Approaches Examined During the Development of BRO
Methods Goup Specific Method Source
Exploration DAC (Nauman & Cygan, 2023)
Value Regularization CDQL (removed) (Fujimoto et al., 2018)
N-Step Returns (Sutton & Barto, 2018)
Network Regularization LayerNorm (Ba et al., 2016)
Weight Decay (Loshchilov & Hutter, 2017)
Spectral Norm (Miyato et al., 2018)
Scheduling N-Step Schedule (Kearns & Singh, 2000)
Discount Schedule (François-Lavet et al., 2015)
Pessimism Schedule
Entropy Schedule
Learning Rate Schedule (Andrychowicz et al., 2021)
Distributional RL HL Gauss (Imani & White, 2018)
Categorical (Bellemare et al., 2017)
Quantile (Dabney et al., 2018)
Learning Replay Ratio (Nikishin et al., 2022; D’Oro et al., 2022)
Appendix C Further Analysis
C.1 Batch Sizes
Ablation of the minibatch size impact on BRO and BRO (Fast) performance across different benchmarks is depicted in Figure 10. The figure shows that using half or even a quarter of the original minibatch size ($256$) does not significantly hurt BRO’s performance.
Figure 10: Performance of BRO and BRO (Fast) with different minibatch sizes across various benchmarks: D&H (Dogs and Humanoid), DMC (all tested DeepMind Control environments), MW (MetaWorld), and MS (MyoSuite).
C.2 Various architecture designs
We examine various architectural blueprints on $5$ DMC and $5$ MetaWorld environments (see Table 3), each with over $10$ seeds per task.
Figure 11: Comparison of architectural backbones: The left diagram shows the backbone of the architecture proposed in this article, while the right diagram displays the backbone of the architecture proposed in (Bjorck et al., 2021). In both models, increasing depth is achieved by adding new residual blocks.
Our starting point was the transformer-based design by Bjorck et al. (2021), termed Spectral. This architecture incorporates recent transformer advancements, moving Layer Norm to the beginning of the residual block to prevent vanishing gradients in deep networks.
As shown in Figure 12, while Spectral performs well overall, its performance on the DMC benchmark, particularly the Dog environment, is weaker. This aligns with findings from Nauman et al. (2024), indicating that Layer Norm is crucial for stability in such complex tasks.
To address this, we replaced spectral norms with Layer Norms in the residual blocks, resulting in the BRO wo first LN architecture (Figure 11). This modification improves performance but still lags behind the full BRO architecture.
Our final experiment examines a simple MLP architecture with Layer Norm before each activation function. Since BRO consists of two residual blocks, we compare it with a 5-layer model, (Dense + LN) x 5. Figure 12 shows that Layer Norm after each Dense layer is effective, and in aggregated IQM, this model is comparable to BRO. However, skip connections in BRO are beneficial for managing complex environments like Dog. In conclusion, BroNet architecture uses Layer Norm and residual blocks for superior robustness and performance in challenging tasks.
Figure 12: Comparison of five architecture designs across different environments: The top plot shows results on 5 DMC and 5 MetaWorld environments, the middle plot focuses on the 5 DMC environments, and the bottom plot highlights the Dog Trot environment. BRO and Spectral architectures each consist of 2 residual blocks, as depicted in Figure 11 (left for BRO and right for Spectral). (Dense + LN) x 5 represents standard MLP networks with 5 hidden layers, each incorporating Layer Norm before activation. Lastly, BRO wo first LN refers to the BRO architecture without Layer Norm in the first Dense block, before the residual connection.
Appendix D Additional Scaling and Time Results
In this section, we provide more details about the time benchmark. The execution time was measured for all agents for each of the $40$ tasks averaged over $2$ random seeds. We ran each agent for $105k$ steps with initial $5k$ exploration steps, $100k$ training steps, and $1$ evaluation.
Experiments were conducted on an NVIDIA A100 GPU with 10GB of RAM and 8 CPU cores of AMD EPYC 7742 processor. All tasks were run separately so the agents could use all resources independently.
We benchmarked all $25$ variants of BRO with $5$ different model sizes and $5$ values of replay ratio. Figure 13 shows how BRO achieves superior performance at lower execution time compared to popular baselines. BRO (Fast) is an attractive alternative, strongly surpassing TD3, CrossQ, and SAC in comparable time, whereas BRO outperforms SR-SAC and TD-MPC2 and, therefore, is a great alternative.
Figure 13: Scatterplot of the performance of agents plotted against normalized execution time. $25$ variants of the BRO with different model sizes and replay ratios dominate the baselines on $40$ tasks in terms of performance and time.
Thanks to different variants of BRO, users have a versatile choice in terms of memory, time, and performance and can use the best agents depending on current constraints. In order to decide between memory, time, and performance tradeoffs, we provide more visualizations of different variants of BRO in Figure 14.
Figure 14: Heatmaps of mean execution time and IQM performance across $40$ tasks of $25$ variants of BRO with various model sizes and replay ratio values. Black lines connect the same interpolated values.
As can be seen in Figure 14 and stated previously, increasing the model size and replay ratio both improve the performance. However, the former is much more efficient in terms of execution time thanks to efficient GPU parallelism. For example, the largest BRO variant ($26.31M$ parameters) with replay ratio $5$ has almost the same execution time as the smallest model ($0.55M$ parameters) with replay ratio $15$, but the performance is much greater.
In some scenarios where GPU memory is limited, good performance can still be achieved by scaling the replay ratio with a time tradeoff.
Table 5: Description of the considered model sizes.
Size Number of block Hidden Size
$0.55M$ 1 BroNet block hidden size of $128$
$1.05M$ 1 BroNet block hidden size of $256$
$2.83M$ 1 BroNet block hidden size of $512$
$4.92M$ 2 BroNet blocks hidden size of $512$
$26.31M$ 3 BroNet blocks hidden size of $1024$
Appendix E Hyperparameters

### v2

@@ 5 Limitations and Future Work
BRO’s larger model size compared to traditional baselines like SAC or TD3 results in higher memory requirements, potentially posing challenges for real-time inference in high-frequency control tasks. Future research could explore techniques such as quantization or distillation to improve inference speed. While BRO is designed for continuous control problems, its effectiveness in discrete settings remains unexplored. Further investigation is needed to assess the applicability and performance of BRO’s components in discrete action MDPs. Additionally, our experimentation primarily focuses on continuous control tasks using proprioceptive state representations. Future research is needed to investigate the tradeoff between scaling the critic and the state encoder in image-based RL.
@@ 6 Conclusions
Our study underscores the efficacy of scaling a regularized critic model in conjunction with existing algorithmic enhancements, resulting in sample-efficient methods for continuous-action RL. The proposed BRO algorithm achieves markedly superior performance within 1 million environment steps compared to the state-of-the-art model-based TD-MPC2 and other model-free baselines. Notably, it achieves over 90% success rates in MetaWorld and MyoSuite benchmarks, as well as over 85% of maximal returns in the DeepMind Control Suite, and near-optimal policies in the challenging Dog and Humanoid locomotion tasks. While some tasks remain unsolved within 1 million environment steps, our findings underscore the need for new standardized benchmarks focusing on sample efficiency to drive consistent progress in the field. The BRO algorithm establishes a new standard for sample efficiency, providing a solid foundation for future research to build upon and develop even more robust RL algorithms.
Acknowledgements
This research was partially supported by the National Science Centre, Poland, under grant nos. 2020/39/B/ST6/01511 and 2023/51/D/ST6/01609, as well as by the Warsaw University of Technology and the University of Warsaw through the Excellence Initiative: Research University (IDUB) program. We also gratefully acknowledge the Polish high-performance computing infrastructure, PLGrid (HPC Center: ACK Cyfronet AGH), for providing computational resources and support under grant no. PLG/2024/017159. Marek Cygan was partially supported by an NCBiR grant POIR.01.01.01-00-0433/20. Piotr Miłoś research was supported by the National Science Center (Poland) grant number 2019/35/O/ST6/03464.
@@ Appendix A Background
with $a^{\prime}\sim\pi_{\phi}$. In this setup, $Q_{\theta}$ is the critic network, $Q_{\bar{\theta}}$ is the value target network, and $\mathcal{D}$ is the replay buffer (Mnih et al., 2015). $Q_{\theta}$ is trained to approximate the Q-value under the policy from which the bootstrap is sampled (Van Seijen et al., 2009; Sutton & Barto, 2018). The policy parameters $\phi$ are updated to seek locally optimal values approximated by the critic (Ciosek & Whiteson, 2020) according to:
@@
$\phi=\arg\max_{\phi}~{}\mathrm{E}_{s\sim\mathcal{D}}~{}\left(Q_{\theta}(s,a)-\alpha\log\pi_{\phi}(a|s)\right),~{}~{}~{}\text{with}~{}~{}~{}a\sim\pi_{\phi}.$ (6)
Appendix B BRO additional details
B.1 Base agent
@@
BRO uses the well-established Soft Actor-Critic (SAC) (Haarnoja et al., 2018) as its base. SAC is a stochastic policy, maximum entropy algorithm (see Eq. 1) that employs online entropy temperature adjustment and an ensemble of two critic networks. SAC models the policy via a Tanh-transformed Gaussian distribution whose parameters are modeled by the actor network.
B.2 Architecture details
@@
In the proposed architecture, we adopt the transformer feedforward blueprint from Vaswani et al. (2017) with a novel layer normalization configuration, as shown in Figure 5. Dropout is omitted. All Dense layers in the BRO have a default width of $512$ units, with a linear layer at both the input and output stages. To increase the model’s depth, we add new residual blocks exclusively. While a similar transformer backbone has been used in previous work (Bjorck et al., 2021), our design choices, detailed in Section E, led us to use layer normalization instead of spectral normalization.
B.3 Scaling
@@
In Figure 3(c), we examine the scaling capabilities of SAC and BRO agents using a vanilla dense network (Fujimoto et al., 2018; Haarnoja et al., 2018; Moskovitz et al., 2021), a spectral normalization ResNet (Bjorck et al., 2021; Cetin & Celiktutan, 2023), and a layer normalization ResNet inspired by previous work (Nauman et al., 2024). As shown in Figure 3(c), increasing the model capacity of a vanilla-dense agent can lead to performance degradation beyond a certain model size. However, for the regularized architectures, we observe behavior similar to the empirically observed scaling properties of supervised models, where increasing model size leads to diminishing returns in performance improvements. Furthermore, we find that the layer normalization ResNet achieves better scaling properties than the spectral normalization architecture. Interestingly, the BRO agent consistently outperforms the baseline SAC across all architectures and network sizes, suggesting an interaction between parameter scaling and other algorithmic design choices. The highest performing SAC agent achieves around $25\%$ of maximal performance, whereas our proposed BRO agent achieves more than $90\%$. Given that the BRO agent performs similarly at $4.92$ million and $26.31$ million parameters, we use the smaller model to reduce the computational burden.
B.4 Scaling replay ratio
@@
B.5 Batch Size
Inspired by recent findings that reducing the batch size can result in significant performance gains for discrete action RL (Obando Ceron et al., 2024), we reduce the number of transitions used for gradient calculation from $256$ to $128$. As shown in Figures 9 & 15, this batch size reduction leads to a marginal improvement in aggregate performance and decreased memory requirements of the algorithm. Interestingly, we find that batch size significantly affects performance, with no single value performing best across all tasks.
B.6 Risk-Neutral Temporal Difference
@@
Recent studies have shown that techniques like layer normalization or full-parameter resets can be more effective at combating overestimation than pessimistic Q-value approximation (Nauman et al., 2024). Since our critic architecture leverages multiple regularization techniques, we disable CDQ and use the ensemble mean instead of the minimum to calculate the targets for actor and critic updates. In Figures 7 & 9, we compare the performance of the baseline BRO to a BRO agent that uses CDQ. Our findings indicate that using risk-neutral Q-value approximation in the presence of network regularization unlocks significant performance improvements without increasing value overestimation.
B.7 Optimistic Exploration
Optimism is an algorithmic design principle that balances exploration and exploitation (Ciosek et al., 2019; Moskovitz et al., 2021). The dual actor setup (Nauman & Cygan, 2023a, b) employs separate policies for exploration and temporal difference updates, with the exploration policy pursuing an optimistic upper-bound Q-value approximation. This approach has been shown to improve the sample efficiency of SAC-based agents (Nauman & Cygan, 2023a). We implement the optimistic policy such that the Q-value upper bound is calculated based on the epistemic uncertainty estimated via the quantile critic ensemble (Moskovitz et al., 2021). Figure 7 shows that using a dual policy setup leads to performance improvements. We observe that these results are particularly pronounced in the early training stages and for smaller networks (Figure 9).
@@ B.7 Optimistic Exploration
7:     Calculate critic target value without CDQ $Q^{k}_{\bar{\theta},\mu}(s^{\prime},a^{\prime})=\frac{1}{2}(Q^{k}_{\bar{\theta},1}(s^{\prime},a^{\prime})+Q^{k}_{\bar{\theta},2}(s^{\prime},a^{\prime}))~{}~{}~{}~{}\text{with}~{}~{}~{}~{}a^{\prime}\sim\pi^{p}_{\phi}(s^{\prime})$
8:     Update critic using pessimistic actor $\theta\leftarrow\theta-\nabla_{\theta}Huber\bigl{(}Q_{\theta}(s,a),(r+\gamma Q^{\mu}_{\bar{\theta}}(s^{\prime},a^{\prime})-\alpha\log\pi^{p}_{\phi}(a^{\prime}|s^{\prime})\bigr{)}$
9:     Calculate pessimistic actor value without CDQ $Q^{\mu}_{\theta}(s,a)=\frac{1}{2K}\sum_{i=1}^{K}(Q^{k}_{\theta,1}(s,a)+Q^{k}_{\theta,2}(s,a))~{}~{}~{}~{}\text{with}~{}~{}~{}~{}a\sim\pi^{p}_{\phi}(s)$
@@
Algorithm 2 BRO training step with changes with respect to standard SAC colored red.
B.8 Approaches Examined During the Development of BRO
Methods Group Specific Method Source
Exploration DAC (Nauman & Cygan, 2023a)
Value Regularization CDQL (removed) (Fujimoto et al., 2018)
N-Step Returns (Sutton & Barto, 2018)
Network Regularization LayerNorm (Ba et al., 2016)
Weight Decay (Loshchilov & Hutter, 2017)
Spectral Norm (Miyato et al., 2018)
Scheduling N-Step Schedule (Kearns & Singh, 2000)
Discount Schedule (François-Lavet et al., 2015)
Pessimism Schedule
Entropy Schedule
Learning Rate Schedule (Andrychowicz et al., 2021)
Distributional RL HL Gauss (Imani & White, 2018)
Categorical (Bellemare et al., 2017)
Quantile (Dabney et al., 2018)
Learning Replay Ratio (Nikishin et al., 2022; D’Oro et al., 2022)
Appendix C Tested Environments
We tested BRO on a variety of $40$ tasks from DeepMind Control Suite (Tassa et al., 2018), MyoSuite (Caggiano et al., 2022) and MetaWorld (Yu et al., 2020). Selected tasks cover various challenges, from simple to hard, in locomotion and manipulation. Table 4 presents the environments with specified dimensions of states and actions. BRO is a versatile agent that can successfully perform tasks of different difficulty and various action and state spaces. Our selection of $40$ tasks focuses on the most challenging tasks from the DeepMind Control Suite (DMC), MetaWorld (MW), and MyoSuite (MS) benchmarks, as identified in previous studies (Hansen et al., 2023). We chose these hard tasks because many easy tasks from these benchmarks can be solved by modern algorithms within 100k environment steps (Hansen et al., 2023; Wang et al., 2024). In the MetaWorld environment, we follow the TD-MPC2 evaluation protocol (Hansen et al., 2023). As such, the environment issues a truncate signal after 200 environment steps, after which we assess if the agent achieved goal success within the 200th step. We do not implement any changes to how goals are defined in the original MetaWorld and we use V2 environments.
Task Observation dimension Action dimension
DeepMind Control
Acrobot-Swingup $6$ $1$
Dog-Trot $223$ $38$
Hopper-Hop $15$ $4$
Humanoid-Run $67$ $24$
Humanoid-Walk $67$ $24$
MetaWorld
Assembly $39$ $4$
Coffee-Push $39$ $4$
Hand-Insert $39$ $4$
Push $39$ $4$
Stick-Pull $39$ $4$
@@ Appendix C Tested Environments
Appendix D Hyperparameters